OpenStudy (anonymous):

An object is thrown upwards and after 2 seconds is 78 m above where it was initially thrown. Assume no air resistance, and determine the initial speed the object was thrown at.

OpenStudy (asnaseer):

you can use this equation of motion to solve this:\[y=ut+{1\over2}at^2\]where: y = distance travelled u = initial velocity a = acceleration (which, in this case, should be equal to -g) t = time

OpenStudy (anonymous):

we are looking for u right?

OpenStudy (asnaseer):

yes

OpenStudy (anonymous):

-78m=u(2) + 1/2 (-9.8)(2)^2 right?

OpenStudy (anonymous):

srry not -78

OpenStudy (asnaseer):

78=...

OpenStudy (asnaseer):

If you are learning about projectile motion, then you might find this helpful: http://hyperphysics.phy-astr.gsu.edu/hbase/mot.html

OpenStudy (anonymous):

i got some weird answer. i got -1372.8

OpenStudy (anonymous):

and thank you soo much for helping

OpenStudy (asnaseer):

try and work out each part step-by-step. so, first what is the value of:\[t^2\]when t=2?

OpenStudy (anonymous):

y = 78m t = 2 a = -9.8

OpenStudy (asnaseer):

yes, but your final answer is wrong, so I'm trying to take you through it one step at a time to se where you may have made a mistake.

OpenStudy (anonymous):

i think i am making a mistake in the re-arranging the equation

OpenStudy (asnaseer):

ok, so first try rearranging the original equation:\[y=ut+{1\over2}at^2\]to get an expression for u. i.e. BEFORE plugging in the values.

OpenStudy (anonymous):

y/u= t + 1/2 at^2

OpenStudy (asnaseer):

no - that is wrong

OpenStudy (anonymous):

then isolate the u? soooo sorry im very terrible at this

OpenStudy (asnaseer):

think of it as:\[A=B+C\]how would you rearrange this to get an expression for B?

OpenStudy (asnaseer):

you need to subtract C from both sides

OpenStudy (anonymous):

a-c

OpenStudy (asnaseer):

yes, therefore:\[y=ut+{1\over2}at^2\]first becomes:\[y-{1\over2}at^2=ut\]

OpenStudy (asnaseer):

next you need to divide both sides by 't'

OpenStudy (anonymous):

OKAYYY!! i was trying to get the easier way out by dividing only. THANK YOU

OpenStudy (asnaseer):

np - do you still want me to check your final answer?

OpenStudy (anonymous):

yup please. 1 min

OpenStudy (anonymous):

48.8 m/s? :s

OpenStudy (asnaseer):

yes - you've got it right - well done!

OpenStudy (anonymous):

LOL THANK YOU SOOO MUCH!!! YOU EXPLAIN VERY GOOD

OpenStudy (asnaseer):

yw - I'm glad I was able to help :)